what is the equation of suppose the length of the board is 7 ft.longer than its widht.what equation would represent the given situation.?
Answer by stanbon(75887)

what is the equation?
Suppose the length of a board is 7 ft.longer than its width.
---------------------------------
What equation would represent the given situation.?
Comment: That depends on the "situation".
If you want area, A = width*length = x(x+7) sq units
----
If you want perimeter, P = 2(width+length+ = 2(x+x+7) = 4x+14 units
----------
Cheers,
Stan H.
----------